OpenStudy (anonymous):

Solve the quadratic equation by factoring. 36x^2 − 1 = 0

OpenStudy (ybarrap):

36x^2 - 1 (6x + 1)(6x - 1)=0 6x + 1 = 0 => x = - 1/6 6x - 1 = 0 => x = 1/6 Ans. x = -1/6 and x = 1/6